### v1

# ObjectClear: Complete Object Removal via Object-Effect Attention

Abstract
Object removal requires eliminating not only the target object but also its effects, such as shadows and reflections. However, diffusion-based inpainting methods often produce artifacts, hallucinate content, alter background, and struggle to remove object effects accurately. To address this challenge, we introduce a new dataset for OBject-Effect Removal, named OBER, which provides paired images with and without object effects, along with precise masks for both objects and their associated visual artifacts. The dataset comprises high-quality captured and simulated data, covering diverse object categories and complex multi-object scenes. Building on OBER, we propose a novel framework, ObjectClear, which incorporates an object-effect attention mechanism to guide the model toward the foreground removal regions by learning attention masks, effectively decoupling foreground removal from background reconstruction. Furthermore, the predicted attention map enables an attention-guided fusion strategy during inference, greatly preserving background details. Extensive experiments demonstrate that ObjectClear outperforms existing methods, achieving improved object-effect removal quality and background fidelity, especially in complex scenarios.
footnotetext: Corresponding author.
1 Introduction
Figure 1: Object Removal Comparison. Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and typically fail to remove associated effects such as shadows and reflections. In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
Recent advances in generative models [27, 24, 21], such as GPT-4o, have shown strong capabilities in image editing and are widely adopted in real-world applications. Among these tasks, object removal has become a key task, allowing users to erase unwanted content from images. However, seamlessly removing both the object and its effects (e.g., shadows and reflections) while preserving the background remains a challenging problem.
While diffusion-based methods [1, 43, 12, 15, 31, 7, 10, 5] have advanced object removal, they still struggle with high-fidelity results. As shown in Fig. 1, existing methods often leave residual artifacts, hallucinate unwanted content, alter backgrounds details, and fail to remove object effects. This is partly because the input mask fully occludes the target object, preventing models from perceiving its surrounding context and associated effects. In addition, the lack of large-scale and public datasets for object-effect removal also limits further progress in the field.
Existing object removal datasets can be categorized into simulated and camera-captured data. (1) Simulated data. These datasets are often constructed by copy-pasting objects [11, 14] or using pretrained inpainting models to generate pseudo ground truth [32]. While this allows low-cost generation of large-scale data, such datasets typically lack object effects such as shadows and reflections, causing models trained on them to struggle with effect removal. (2) Camera-captured data. Some works [29, 33] leverage fixed-viewpoint videos to extract paired frames, but this constrains the foreground types to moving objects. Others [34, 39, 36] capture image pairs before and after object removal, but these datasets are often not publicly available and small in scale due to the high costs.
To bridge this gap, we introduce OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects. For the camera-captured data, we annotate precise object and object-effect masks, which serve as critical supervision during training. For the simulated data, we leverage these masks to compute accurate alpha maps, enabling realistic alpha blending of RGBA object layers with effects and high-quality backgrounds. We further extend the simulation pipeline to include multi-object scenarios, enhancing model robustness in challenging cases involving occlusion and object interactions, as shown in the last two rows of Fig. 5. By combining the realism of captured data with the scalability of simulation, OBER offers a high-quality, large-scale, and diverse dataset with a total of 12,715 training samples. In addition, we introduce two new benchmarks: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth), to support future research.
Built upon OBER, we propose a novel framework, ObjectClear, which introduces an Object-Effect Attention (OEA) mechanism that adaptively focuses on the foreground removal region, including both the object and its associated effects, by supervising cross-attention maps with object-effect masks. The OEA module effectively decouples foreground removal from background reconstruction, enabling more precise and complete object elimination. In addition, the predicted attention map supports an Attention-Guided Fusion (AGF) strategy during inference, helping to preserve background details, as shown in Fig. 7. This design enables the model to localize editable regions adaptively and offers a potential solution to the long-standing fidelity issues in diffusion-based image editing.
Our contributions are summarized as follows:
  • We propose OBER, a high-quality and large-scale hybrid dataset for object removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex multi-object scenarios across both simulated and camera-captured settings.
  • We introduce a novel framework, ObjectClear, which incorporates an Object-Effect Attention mechanism that adaptively focuses on foreground removal regions and enables an Attention-Guided Fusion strategy, improving object removal quality and background fidelity.
  • We present two new benchmarks, OBER-Test and OBER-Wild, providing a foundation for future research. Our method achieves superior performance on all benchmarks, outperforming existing approaches in terms of both quantitative metrics and visual quality.
2 Related Work
Image Inpainting. Image inpainting is a long-standing visual editing task that aims to seamlessly reconstruct pixels within a given mask. Early approaches predominantly adopt generative adversarial networks (GANs) [8], but often suffer from limited realism and diversity [17, 22, 26, 40]. With the rapid advances in diffusion models [28, 30, 9, 24], many methods [2, 20, 21, 43, 12, 38] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting. In this work, we adapt the SDXL-Inpainting model [1] for photorealistic completion. However, despite their strong generative capabilities, existing inpainting models often lack awareness of object-induced effects (e.g., shadows and reflections), leading to incomplete or visually inconsistent object removal results.
Object Removal. Object removal is a specialized form of image inpainting that requires accounting for object effects to ensure complete and realistic removal, which diffusion-based models [34, 11, 19, 39, 15, 4, 43, 7, 5, 36, 10, 31, 33] currently dominate. A common strategy involves curating high-quality triplet datasets. For instance, ObjectDrop [34] builds a real-world dataset by having photographers capture the same scene before and after the removal of a single object. However, its limited scale and lack of public availability hinder broader adoption. To address data scalability, methods like SmartEraser [11] and Erase Diffusion [19] rely on synthetic datasets using segmentation or matting techniques to extract foreground objects, but these typically lack annotations of object effects, limiting the models’ ability to remove shadows or reflections. To enhance realism, methods such as LayerDecomp [36] and OmniPaint [39] create costly camera-captured data. OmniPaint [39] takes one step further to annotate unlabeled images by a model trained on small-scale real data, while RORem [15] involves human annotators to ensure the quality of annotated data. In parallel, works like RORD [29] and OmniEraser [33] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames with and without target objects while preserving natural object effects. To eliminate reliance on curated datasets altogether, other methods opt for test-time optimization [5, 31, 10] or providing plug-and-play solutions [7]. However, all these models rely on implicit learning of object effects without explicitly modeling effect maps, making it difficult to ensure that no unnecessary background changes occur.
3 Methodology
To achieve accurate removal of target objects along with their associated visual effects, such as shadow and reflection, we propose a comprehensive framework consisting of a data curation pipeline (OBER Dataset) and a dedicated object removal model (ObjectClear). These two components are detailed in Sec. 3.1 and Sec. 3.2, respectively.
3.1 OBER Dataset
ObjectDrop [34] shows that even small-scale datasets, when captured with careful camera control over scene consistency, can significantly enhance model generalizability in object removal tasks. However, due to the lack of public access to such datasets and the high cost of real-world data collection, subsequent works often face limitations in dataset scale. To address this challenge, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability. As illustrated in Fig. 2, OBER is a hybrid dataset consisting of two parts: (1) a small set of camera-captured images adhering to physical realism, and (2) a larger set of simulated images generated by compositing foreground objects, extracted from the real data in (1), onto diverse background scenes.
Figure 2: Dataset Construction Pipeline of OBER. The dataset combines both camera-captured and simulated data, featuring diverse foreground objects and background scenes. It provides rich annotations, including object masks, object-effect masks, transparent RGBA object layers, and complex multi-object scenarios for training and evaluation.
Camera-Captured Data. (1) Capture Paired Images. Following the approach of ObjectDrop [34], we use fixed cameras to construct a counterfactual dataset by capturing each scene before and after the removal of a single object, while keeping all the other factors unchanged. For each pair, the image with the object is used as the input $I_{in}$, and the image without the object serves as the ground truth $I_{gt}$. In total, we collected 2,875 such counterfactual pairs with 2,715 for training and 160 for testing. Our dataset encompasses a wide variety of everyday objects commonly found in indoor (e.g., benches, cups, dolls) and outdoor scenes (e.g., pedestrians, vehicles). These image pairs also preserve visual effects such as shadows and reflections. To reduce potential pixel-level misalignment, we downsample and crop all images to a fixed training resolution (i.e., 512 $\times$ 512). (2) Segment Object Mask $M_{o}$. To obtain the object mask $M_{o}$ as network input alongside the input image $I_{in}$, we apply off-the-shelf detection and segmentation models, such as DINO [41] and SAM [13] to $I_{in}$. (3) Segment Object-Effect Mask $M_{o}\cup M_{e}$. Different from previous methods, we propose to explicitly model the object effects. Therefore, we introduce an object-effect mask $M_{fg}$ that covers both the object $M_{o}$ and its associated visual effects $M_{e}$ in our dataset. Unlike the coarse object-effect masks provided in the RORD dataset [29], we construct $M_{fg}$ in an efficient and accurate way by computing the pixel-wise difference between $I_{in}$ and $I_{gt}$. Pixels with differences above a predefined threshold are regarded as part of the object-effect mask. This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its effects.
Simulated Data. With the high-quality real data collected, we further scale up the dataset with a carefully designed simulation pipeline. (1) Collect Background Images. We begin by downloading high-quality background images from the Internet, which offer a diverse range of background scenes. To select backgrounds with flat regions suitable for object placement, we first apply Mask2Former [6] to segment flat areas corresponding to semantic classes such as “road, sidewalk, grass, floor”, as illustrated in Fig. 2(b). We then refine the selection by computing the gradient of the depth map generated by Depth Anything V2 [37], filtering out regions with significant depth variation to ensure that the inserted objects are placed on visually flat surfaces. (2) Collect Foreground Objects with Effects. Based on the camera-captured paired data $(I_{in},I_{gt})$, along with the object mask $M_{o}$ and effect mask $M_{e}$, we compute the alpha map of the foreground object with effects $I_{oe}$ using Eq. 1, where $\varepsilon$ is a small constant added to avoid division by zero. For subsequent compositing, we manually categorize the foreground objects into eight groups based on their shadow direction (Fig. 2(b)).
$\alpha(p)=\begin{cases}0,&\text{if pixel }p\in\text{background area $\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if pixel }p\in\text{object area $M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if pixel }p\in\text{effect area $M_{e}$}\\ \end{cases}$ (1)
(3) Blend Objects with Backgrounds. We randomly sample a background image $I_{bg}$ and a foreground object image with effects $I_{oe}$, and synthesize a composite image using alpha blending: $I_{comp}=(1-\alpha)\cdot I_{bg}+\alpha\cdot I_{oe}$. Beyond single-object cases, we also synthesize multi-object data by compositing multiple foreground objects with the same lighting directions, thereby covering scenarios involving object occlusions, as illustrated in Fig. 2(b). In total, we generate 10,000 composite images, significantly enriching the diversity and scalability of our dataset. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness and background preservation, as discussed and compared in Table 2.
3.2 ObjectClear
Figure 3: The Framework of ObjectClear. Given an input image and a target object mask, ObjectClear employs an Object-Effect Attention mechanism to guide the model toward foreground removal regions by learning attention masks. The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
Our ObjectClear is built upon SDXL-Inpainting [1]. While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $<z_{t},I_{in},M_{o},c>$. This design encourages the model to better attend to the effect of the target object by leveraging its visual features. Moreover, it facilitates more effective utilization of background information behind the object in cases involving moving transparent objects, such as glass cups. Apart from the input design, we further introduce an Object-Effect Attention Mechanism and an Attention-Guided Fusion Strategy to explicitly attend object-effect regions and enhance background preservation. Detailed descriptions of these components are provided below.
Object-Effect Attention. To enable the model to better attend to both the target object region and its associated visual effect regions, we follow the approach of Li et al. [16] by stacking text and visual object information as guidance for the cross-attention layers in the base model. Specifically, the text prompt is expressed as “remove the instance of”, and the visual object is obtained by applying a dot product between the input image $I_{in}$ and $M_{o}$, as illustrated in Fig. 3(a). These two modalities are then encoded into text embeddings and visual embeddings using the CLIP [25] text and vision encoders, respectively. To unify their representation spaces, the visual embeddings are further projected into the same dimensional space as the text embeddings using a lightweight trainable MLP composed of two linear layers. The resulting text embeddings and projected object embeddings are then stacked and used as guidance in the cross-attention blocks of the base model. To encourage the model to focus more accurately on the object and its effect regions, we introduce a mask loss $\mathcal{L}_{mask}$. Concretely, we extract the cross-attention maps corresponding to the visual embedding tokens and denote them as $\mathbf{A}$, which we supervise with the annotated foreground object-effect masks $M_{fg}$ from our OBER dataset. Following [35, 23], $\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in the foreground. This objective can be formulated as:
$\mathcal{L}_{mask}=\mathrm{mean}({\mathbf{A}}[1-M_{fg}])-\mathrm{mean}({\mathbf{A}}[M_{fg}])\text{,}$ (2)
where []delimited-[][$\cdot$][ ⋅ ] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
Attention-Guided Fusion. Interestingly, we observe that applying Object-Effect Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects. Based on this observation, we propose Attention-Guided Fusion strategy that utilizes attention maps during inference to seamlessly combine the generated result with the original image. Specifically, we extract the first-layer cross-attention map corresponding to the object embedding during inference. This attention map is then upsampled to match the spatial resolution of the original image, forming a soft estimate of the object-effect region. To enhance the spatial smoothness and avoid hard edges, we apply a Gaussian blur to the upsampled attention map, producing a softened object-effect mask. This blurred mask is then used to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image. This strategy significantly reduces undesired background changes caused by the diffusion process and VAE reconstruction, while preserving high-fidelity object removal. Unlike BrushNet [12], which relies on user-provided object masks to guide blending, our approach adaptively produces the object-effect mask from our Object-Affect Attention.
4 Experiments
Table 1: Quantitative Comparisons with State-of-the-Art Methods. The best and second performances are marked in red and orange, respectively. Although ObjectClear takes only the object mask as input, it outperforms previous methods, even when those methods are provided with masks that cover both the object and its associated effect regions.
Mask Datasets RORD-Val OBER-Test OBER-Wild
[…]
Object-Effect Mask SDXL-INP [1] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.8356
PowerPaint [43] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8878
BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.6415
DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.9150
CLIPAway [7] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.8646
FreeCompose [5] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.8976
Attentive Eraser [31] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.9095
RORem [15] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.9157
Object Mask SDXL-INP [1] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.8044
PowerPaint [43] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.8538
BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.7296
DesignEdit [7] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.9002
CLIPAway [10] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.8646
FreeCompose [5] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.8688
Attentive Eraser [31] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.9008
RORem [15] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.9025
ObjectClear (Ours) 26.40 30.75 0.0802 0.0350 32.98 35.14 0.0352 0.0099 0.9172
Implementation. Our proposed method is built on SDXL-Inpainting [1] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$. Training is conducted with a total batch size of $32$ on 8 A100 GPUs for 100k iterations and a learning rate of 1e-5. All experiment results are attained with a guidance scale of $7.5$ and $20$ denoising steps.
Evaluation Data. We evaluate the object removal performance of our method, ObjectClear, against existing approaches on three test datasets: (1) RORD-Val: RORD [29] is a widely used object removal dataset. Each sample contains an image pair (images with and without the target object) and a coarse mask covering both the target object and its associated effect. To avoid duplicated scenarios, we randomly select one image per scene, resulting in a RORD-Val with 343 samples. Since most of the object removal methods, including ObjectClear, require only the object mask as input, we augment RORD-Val with accurate object masks manually to enable a more comprehensive evaluation. (2) OBER-Test: As described in Sec. 3.1, we split our collected OBER dataset into a training set and a test set. The test set contains 160 samples, each comprising image pairs along with precise masks for both the object and its effect. (3) OBER-Wild: To further assess the performance of different methods on out-of-distribution scenarios in the wild, we collect 300 high-quality images featuring objects with associated effects (shadows or reflections) from the Internet. We annotate the object masks using DINO [41] and SAM [13], and manually annotate the effect masks. Note that the OBER-Wild set does not include the removal ground truth.
Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects.
Figure 5: Object Removal on OBER-Wild. Our ObjectClear not only effectively removes single objects with their shadows and reflections (top five samples), but also accurately removes the target object when multiple mutually occluding objects exist (bottom two samples).
Evaluation Metrics. For RORD-Val and OBER-Test, where ground truth is available, we evaluate performance using fidelity metrics: PSNR and PSNR-BG (computed only on background regions), as well as two perceptual metrics: LPIPS [42] and CLIP [25] (feature distance). For OBER-Wild, which lacks ground truth, we employ ReMOVE [3] to measure the consistency between the object removal region and the surrounding background.
Figure 6: Object-Effect Attention Map. ObjectClear achieves a relatively high recall in the object-effect attention map, with recall values increasing as the denoising step progresses. The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
4.1 Comparisons with State-of-the-Art Methods
We compare ObjectClear with open-source state-of-the-art methods in both image inpainting and object removal. The inpainting methods include SDXL-INP [1], PowerPaint [43], and BrushNet [12]. The object removal methods include CLIP-Away [7], DesignEdit [10], RORem [15], FreeCompose [5], and Attentive Eraser [31]. Following MULAN [32], we use the text prompt “an empty scene” for the text-guided baselines when necessary.
Quantitative Evaluation. Unlike our ObjectClear, which explicitly handles both the target object and its associated effects, most baseline methods operate only within the masked regions, overlooking surrounding areas that are visually affected by the object but not explicitly included in the mask. To ensure a fair and comprehensive comparison, we evaluate all methods under two mask conditions: (1) the input mask covers only the target object, and (2) the input mask covers both the target object and its visual effects. These are referred as Object Mask and Object-Effect Mask settings in Table 1, respectively. As shown in Table 1, ObjectClear achieves state-of-the-art performance across all test sets and metrics. Notably, even when using only the object mask, ObjectClear surpasses methods that rely on both object and effect masks. In particular, it achieves a significant advantage in the PSNR-BG metric, highlighting its superior ability to preserve background consistency with the input.
Qualitative Evaluation. Qualitative results are shown in Fig. 4 and Fig. 5. Generation-based inpainting approaches, such as SDXL-INP [24], BrushNet [12], and PowerPaint [43], often generate new objects within the masked regions while fail to remove the effects of the removed objects. In contrast, previous object removal methods, such as Attentive Eraser [31], DesignEdit [10], and RORem [15] demonstrate strong performance in removing the target object itself, but stil fail to eliminate the associated effects. Thanks to our carefully curated training data and the proposed object-effect attention mechanism, ObjectClear effectively removes not only the target object but also its associated shadows and reflections. Moreover, as shown in the last two rows of Fig. 5, ObjectClear effectively handles complex cases involving multiple mutually occluding objects.
Table 2: Quantitative Results of Ablation Studies. Based on our camera-captured data (CC Data), our object removal performance improves progressively by adding the object-effect mask loss ($\mathcal{L}_{mask}$) and simulated training data (Sim. Data). Further improvement is achieved through the proposed Attention-Guided Fusion (AG Fusion).
[…]
(a) 27.29 27.96 0.0910 0.0247
(b) 27.56 28.37 0.0845 0.0217
(c) 28.00 28.81 0.0818 0.0190
(d) 32.98 35.14 0.0352 0.0099
4.2 Ablation Studies
The Effectiveness of the Object-Effect Mask Loss $\mathcal{L}_{mask}$. The object-effect mask loss is designed to explicitly guide ObjectClear to attend to both the target object and its associated effect regions, leading to more thorough and natural removal results. As shown in the chart on the left of Fig. 6, the predicted object-effect attention map achieves a high recall of 0.97. Besides, the attention map in the last column of Fig. 6 on the right, which is from the final denoising step (Step 20), effectively reveals not only the object’s shadow (marked with a red box), but also the object’s reflection in the mirror (marked with a green box), altogether leading to complete and visually coherent object removal results. From Table 2 (a) to (b), improvements can be seen across all metrics with $\mathcal{L}_{mask}$ added, indicating its effectiveness in improving object removal accuracy.
The Effectiveness of Simulated Data. The simulated data effectively enhances the background diversity and scenario complexity of the training set, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c). Our carefully designed simulation pipeline, especially the multi-object data synthesis, enables ObjectClear to handle challenging cases involving multiple mutually occluding objects, as demonstrated in the last two rows of Fig. 5. More visualized results can be found in the appendix.
Effectiveness of the Attention-Guided Fusion. The attention-guided fusion operation is designed to preserve background consistency during object removal, a common challenge in generation-based methods such as ObjectDrop [34] and RORem [15], as illustrated in Fig. 7. As the denoising process progresses, the proposed object-effect attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of background images. This strategy significantly improves background preservation, as evidenced by notable gains in PSNR-BG. Consequently, we observe a substantial boost in overall performance, as reflected in the quantitative results from (c) to (d) in Table 2, where PSNR improves markedly from $28.00$ to $32.98$.
Figure 7: Effectiveness of Attention-Guided Fusion. The difference maps visualize the pixel-wise differences between the input and output images. RORem [15] fails to remove the reflection of the balloon and introduces changes to the background, resulting in relatively high difference values in the background area and low values in the reflection region. In contrast, both ObjectDrop [34] and our ObjectClear successfully remove the reflection. Notably, ObjectClear better preserves the background, thanks to our predicted object-effect map and proposed attention-guided fusion strategy.
4.3 More Applications
ObjectClear can be flexibly extended to various applications, such as object insertion and movement. We also utilize our OBER dataset to train an insertion network that is able to generate natural object effects. Please refer to the appendix for more details. As shown in Fig. 8, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
Figure 8: Object Insertion and Movement. In addition to accurately inserting or repositioning objects, our ObjectClear also generates plausible and natural shadows and reflections accordingly.
5 Conclusion
We introduce ObjectClear, a practical framework for object removal that ensures high-quality object-effect removal and background-preserving results across diverse real-world scenarios. Our method employs an Object-Effect Attention (OEA) mechanism to adaptively focus on removal regions, including both the object and its effects, coupled with Attention-Guided Fusion (AGF) to preserve background details accurately. We also propose OBER, a large-scale and diverse hybrid dataset combining camera-captured and simulated data. With the carefully designed mechanism and curated dataset, ObjectClear achieves robust and superior performance, effectively addressing existing challenges in object removal tasks. In addition, ObjectClear can be easily adapted to various applications, such as object insertion and object movement, and consistently achieves superior performance. These advancements position ObjectClear as a reliable solution for practical image editing, laying a solid foundation for future research in controllable image processing through local editable regions.
Appendix
In this supplementary material, we provide additional discussions and results to complete the main paper. In Sec. A, we present further details on data augmentation during training and other extended applications. In Sec. B, we provide more details of the proposed OBER dataset, including statistics and some examples for demonstration. In Sec. C, we report additional results, such as more results on ablation studies, a user study, experiments with user strokes, more results of object insertion and movement, and additional comparisons on in-the-wild data.
A More Details of ObjectClear
A.1 Training Augmentations
During training, we apply on-the-fly random cropping to enable the network to learn object removal across varying object sizes. We also apply color augmentation and random flipping to enhance model robustness. To improve the model’s robustness to the estimated or user-provided coarse mask, several previous methods [34, 11, 33] have introduced mask augmentation techniques. In line with these approaches, we also apply dilation and erosion to the input mask during training. Different from previous practice, our method employs an object-aware dilation and erosion strategy, where the dilation and erosion kernel size is adaptively determined based on the size of the object. As demonstrated in the qualitative results in Sec. C.3, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
A.2 More Details for Other Applications
In this section, we further discuss implementation details of other applications, including object insertion and object movement.
Object Insertion. The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear (Fig. 3 of the main paper), thus receiving the input tuple of $<z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for object-effect attention). However, $I_{in}$, $I_{GT}$, and $c$ are constructed in an “reverse” way compared with the removal network.
While the ground truth $I_{GT}$ is the original image with the object accompanied by natural effects, $I_{in}$ is the one with the object simply copied and pasted on the background. Specifically, to generate $I_{in}$, we first extract the object from $I_{GT}$ with the object mask $M_{o}$ only, and then we paste it onto the corresponding background image. In addition to the image pair, the input text $c$ is also reversed to “insert the instance of”. Notably, the object-effect attention map is still supervised by $M_{fg}$. In this context, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
During inference, we directly paste an object (without effects) onto the background scene, along with its object mask $M_{o}$, which forms the input image $I_{in}$. The network then generates the output image where both the object and its generated effects are harmoniously inserted. Since the insertion network also integrates the object-effect attention, we can also apply Attention-Guided Fusion to preserve the background fidelity while generating realistic object effects.
Object Movement. To enable object movement, we combine the ObjectClear and insertion network introduced above. Specifically, we first apply ObjectClear to remove the target object along with its associated effects (e.g., shadows or reflections), resulting in a clean object-free background. The object is then extracted using its given object mask. Users are allowed to specify a new location and optionally adjust the object scale before the object is re-inserted into the clean background. With our insertion network, the object is harmonized with the new context by generating realistic effects. This two-stage approach supports controllable object movement while ensuring visual realism and consistency.
B More Details of OBER Dataset
Table 3: Dataset Overview. The OBER dataset consists of camera-captured and simulated training data, as well as two testing sets: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth). All data subsets are annotated with object masks and object-effect masks.
Properties Training Set Testing Set
Camera-Captured Simulated OBER-Test OBER-Wild
#Image Pairs 2,715 10,000 160 300
w/ Ground Truth ×
w/ Object Mask
w/ Object-Effect Mask
Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the Counterfactual dataset is not publicly available.
Description RORD [29] MULAN [32] DESOBAv2 [18] Counterfactual [34] OBER (ours)
Tasks Object Removal × ×
Effect Removal × ×
Object-Effect Removal × ×
Annotations Object Mask ×
Effect Mask × × ×
Object-Effect Mask ✓(Coarse) × ×
RGBA Objects × × × ×
Multi Objects ×
Effect Types Shadow × Shadow Shadow Shadow & Reflection
Camera-Captured GT × ×
Overview. Table 3 provides an overview of our OBER dataset. The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks. For evaluation, we provide two test subsets: OBER-Test (includes 160 pairs with ground truth) and OBER-Wild (includes 300 pairs without ground truth), where object masks and object-effect masks are also available. In addition, we showcase some samples from our OBER dataset in Fig. 9, including camera-captured data and simulated data, alongside with their annotations such as object masks, object-effect masks, and RGBA object foregrounds.
Reflection Pair Simulation. Thanks to the strong priors of generative models, we observed that the model trained without reflection data was still able to remove some simple reflection effects, a finding consistent with the conclusions drawn in ObjectDrop [34]. However, the model often struggled with more challenging reflection removal cases, such as when the object and its reflection were spatially separated or when the reflection was heavily distorted by surface ripples. To address this limitation, we adopted a human-in-the-loop strategy to collect paired reflection data. Specifically, we used our trained model to perform inference on 200 real-world reflection images and manually selected 50 high-quality results (as shown in Fig. 9(b)), which were then added as an important supplement to the training data. We found that even a small amount of high-quality reflection data pairs could considerably improve the model’s ability to generalize across diverse reflection scenarios.
Comparison with Existing Datasets. We compare our OBER dataset with existing datasets, including those focused on object removal (RORD [29], MULAN [32], Counterfactual Dataset [34]) and one on shadow removal (DESOBA-v2 [18]). As summarized in Table 4, our OBER is the only dataset that includes all three types of mask annotations and RGBA foreground objects, enabling a wide range of tasks such as object removal, effect removal, and joint object-effect removal.
While RORD [29] is among the first to provide object-effect masks, those masks are coarse without separate annotations for objects and their effects. As a result, it cannot support independent removal of either the object or the effect, but only their joint removal. In contrast, our OBER dataset provides precise and separate masks for objects and their effects, enabling more flexible and fine-grained removal tasks. In addition, MULAN [32] focuses solely on object removal without addressing associated effects such as shadows or reflections. DESOBA-v2 [18] targets only effect removal and thus does not handle the objects. Moreover, the ground truth images in both datasets are not captured by cameras but are instead synthesized using generative models, which may limit their realism. The Counterfactual Dataset proposed in ObjectDrop [34] is not publicly available, and it only includes object masks without object-effect masks. It also lacks foreground RGBA data for objects and associated effects, limiting its extensibility and applicability to more advanced object editing tasks.
Figure 9: Samples from our OBER Dataset. We provide some samples from our OBER dataset including both camera-captured data and simulated data. While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec. 3.1 in the main paper), which could be used to construct realistic simulated samples. In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets (Table 4), thanks to our reflection pair simulation strategy detailed in Sec. B.
C More Results
C.1 Results for Ablation Study
Effectiveness of $\mathcal{L}_{mask}$. Thanks to the rich annotations in our OBER dataset, the object-effect mask enables a supervision loss, denoted as $\mathcal{L}_{mask}$, which guides the cross-attention layers to focus on both the object and its associated effects, while preserving background textures. This supervision facilitates a decoupled optimization of object removal and background reconstruction. As shown in Fig. 10, with the mask loss $\mathcal{L}_{mask}$, the network can adaptively identify the object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes). This leads to more accurate and complete removal of objects and their shadow effects, without mistakenly erasing unrelated background content.
Figure 10: Effectiveness of $\mathcal{L}_{mask}$. It could be observed that when training without the $\mathcal{L}_{mask}$ supervision, the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left). In contrast, when supervised with $\mathcal{L}_{mask}$, the attention mask (shown in yellow boxes) could identify the removal regions well, thus leading to more accurate and complete removal results.
Effectiveness of Simulated Data. To balance data realism and scalability, in addition to the camera-captured data, we scale up our OBER dataset with a carefully designed simulation pipeline. Our simulation data is generated by compositing the foreground RGBA object (extracted from camera-captured data) onto diverse backgrounds. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness when mutually occluding objects exist, as shown in Fig. 11 (left). Furthermore, our simulated reflection image pairs (discussed in Sec B) greatly improve the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples, as shown in Fig. 11 (right).
Figure 11: Effectiveness of Simulated Data. Since our simulation data includes multi-object compositions, training with such data enables the model to accurately remove the target object and its associated effects while preserving unrelated object effects (left). In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples (right).
Effectiveness of Attention-Guided Fusion. The attention map supervised by $\mathcal{L}_{mask}$ supports the Attention-Guided Fusion (AG Fusion) strategy during inference. It helps to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image. Such practice effectively reduces undesired background detail changes caused by VAE reconstruction errors and the diffusion process, thereby greatly preserving the background fidelity. In Fig. 12, we visualize the background detail changes by showing the difference maps between the generated image and the corresponding input, where a clear improvement on background preservation could be observed.
Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion). We visualize the background detail changes by showing the difference maps between the generated image and the corresponding input. In ideal cases, the difference maps should show large values only in the object removal regions, including the object and its effects, while showing little difference in the unrelated background. It could be observed that when AG Fusion is not employed, although the difference in object removal regions is significant, the difference in the unrelated background is also noticeable. However, with the AG Fusion strategy, we almost eliminate the undesirable background difference thanks to the accurate attention map predicted.
C.2 User Study
To enable a more comprehensive evaluation, we conducted a user study on object removal results for in-the-wild images. We compared ObjectClear with two representative state-of-the-art methods: PowerPaint [43] and RORem [15]. For a fair comparison, we evaluated each baseline method under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask. In contrast, all ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of the baselines.
We invite a total of 30 participants for this user study. Each volunteer was presented 40 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result. Participants were asked to select the best removal result based on two criteria: the realism of the object region and the preservation of background details.
As summarized in Table 5, ObjectClear outperforms the baselines under both mask settings. Notably, although baseline methods benefited from access to object-effect masks, our ObjectClear still won more user preference with the object mask only.
Table 5: User Study. Our ObjectClear is preferred by human voters over other two representative state-of-the-art methods (PowerPaint [43] and RORem [15]) under both settings. Although the compared methods won more user preference given the object-effect mask instead of the object mask, our ObjectClear still received far more user preference with the object mask only.
Methods PowerPaint (Object Mask) [43] PowerPaint (Object+Effect Mask) [43] ObjectClear (Ours)
User Preference 2.68% 7.03% 90.29%
Methods RORem (Object Mask) [15] RORem (Object+Effect Mask) [15] ObjectClear (Ours)
User Preference 3.51% 11.80% 84.69%
C.3 Results with User Strokes
In practical applications, users often interact with visual systems through imprecise or casually drawn inputs, such as rough scribbles or incomplete masks. These inputs may vary significantly in shape, location, and accuracy. Therefore, it is essential for a robust object removal network to effectively process such arbitrary mask inputs without relying on carefully crafted annotations. Benefiting from our mask augmentation strategy and object-effect attention mechanism, our network demonstrates strong robustness to diverse mask inputs. In this subsection, we simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks, as illustrated in Fig. 13.
Figure 13: Results with User Strokes. We simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks.
C.4 More Comparisons on In-the-wild Data
In this subsection, we compare ObjectClear with state-of-the-art methods, including SDXL-INP [1], PowerPaint [43], Attentive Eraser [31], DesignEdit [10], and RORem [15], on in-the-wild data. To ensure a fair comparison, we consider two settings: (1) All methods use the same object mask; (2) While other methods use finely annotated object-effect masks, our ObjectClear still uses the object mask. The results under these two settings are shown in Fig. 14 and Fig. 15 separately. Under the first setting, when given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts. Under the second setting, while other methods do remove the shadow thanks to the given effect region, they also alter or remove the original background content in an undesirable way. In contrast, our ObjectClear removes the object effect accurately with the object mask only while preserving the background with high fidelity.
Figure 14: Object Removal on OBER-Wild - Condition on Object Mask. Given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts.
Figure 15: Object Removal on OBER-Wild - Comparison between methods conditioned on Object-Effect Mask (others) and Object Mask (ours). Since existing methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated object-effect masks for a fair comparison. Although these methods are able to remove shadows with the additional effect region, they often introduce undesirable changes to the original background. In contrast, ObjectClear effectively removes the object and its associated effects using only the object mask, while preserving the background content with high fidelity.
C.5 Limitations
While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios. Specifically, (1) when multiple shadows heavily overlap, the model struggles to accurately disentangle and localize the effect corresponding to the target object; and (2) in cases involving extremely large and long shadow regions, ObjectClear may fail to completely remove the effect, occasionally leaving behind residual artifacts. Enhancing the disentanglement of complex effect regions and improving the removal of large-scale effects remain important directions for future research.